An individual consumer consumes two commodities X1 & X2. The utility function is U = X1^0.4 X2^0.6. The price of commodity one is P1 = Rs. 3.00, the price of commodity two is P2 = Rs. 4.00, the individual's income per period is Rs. 108. Determine the utility maximizing level of X1 & X2 and derive the fans curve for the two commodities.
1
Expert's answer
2020-11-11T08:27:35-0500

At the utility maximizing level of X1 & X2:

"MU1\/P1 = MU2\/P2" or MU1/MU2 = P1/P2,

"MU1 = U'(X1) = 0.4(X2\/X1)^{0.6} ,"

"MU1 = U'(X1) = 0.6(X1\/X2)^{0.4},"

So, we get:

"\frac{2X2} {3X1} = \frac{ 3} {4} ,"

X1 = 8/9 X2,

According to the budget equation:

3X1 + 4X2 = 108, we get:

"3\u00d78\/9\u00d7X2 + 4X2 = 108,"

"20\/3\u00d7X2 = 108,"

X2 = 16.2 units,

"X1 = 8\/9\u00d716.2 = 14.4" units.